OpenStudy (volleyballlover55):

A cone-shaped paper cup can hold 35 cubic inches of food. If the radius of the cone is 2.4 inches, what is the height of the cone? Use 3.14 for pi. Round your answer to the nearest hundredth. 5.81 inches 6.57 inches 7.89 inches 8.89 inches

OpenStudy (volleyballlover55):

@johnweldon1993

OpenStudy (kittiwitti1):

Okay, let's first focus on the context.\[\text{"can hold"}=Volume_{cone}=35\text{ in}^{3}\]So the volume of the cone is 35 cubic inches... now add that to the formula.\[V_{cone}=\frac{1}{3}\times base\times height\]

OpenStudy (volleyballlover55):

whats the base and height?

OpenStudy (kittiwitti1):

Since the base is a circle, you have to incorporate the area of a circle into this equation which is \(A_{circle}=\pi r^{2}\)

OpenStudy (kittiwitti1):

The problem wants you to use 3.14 instead of \(\pi\), and the radius given is \(r=2.4\text{ in}\). So the area of the circle becomes \(A_{circle}=3.14\times(2.4)^{2}\). Now we need to put this back into the original formula for the cone's volume:\[V_{cone}=\frac{1}{3}\times A_{circle}\times h\]\[V_{cone}=\frac{1}{3}\times(3.14\times(2.4)^{2})\times h\] I set height as a variable, \(h\)

OpenStudy (kittiwitti1):

You have the volume of the cone as 35 cubic inches: \[35=\frac{1}{3}\times3.14(5.76)\times h\]Now all you need to do is isolate the h to find its value.

OpenStudy (kittiwitti1):

@volleyballlover55 everything good? I hope my explanation made sense o.o

OpenStudy (volleyballlover55):

yes it did so whats the final answer?

OpenStudy (kittiwitti1):

Did you try and solve for \(h\)?

OpenStudy (volleyballlover55):

let me try... give me a sec..

OpenStudy (volleyballlover55):

i got 5.8!

OpenStudy (volleyballlover55):

so its A?

OpenStudy (kittiwitti1):

Sorry, I'm working it out

OpenStudy (volleyballlover55):

ok..

OpenStudy (kittiwitti1):

Yup. I got 5.8 as well.\[\text{Good job!}\]

OpenStudy (volleyballlover55):

ok next! A can of chicken stock measures 3.5 inches in diameter and is 5 inches tall. What is the volume of the can? Use 3.14 for pi. Round your answer to the nearest hundredth. 38.50 cubic inches 48.08 cubic inches 92.46 cubic inches 192.33 cubic inches

OpenStudy (kittiwitti1):

Sorry, I've got to study for my final exam ; - ;

OpenStudy (kittiwitti1):

I'll give you a hint though, since you have the formula d = 3.5, so radius is half of that

OpenStudy (kittiwitti1):

put the radius in the "base" section (area of circle) and then calculate the whole thing sorry, gtg! ; - ;